OpenStudy (anonymous):

Will Fan and Medal!!! (√5-x) --------- = ????? (√x+5)

OpenStudy (anonymous):

no0 the answer choices are A. {x|x<5}

OpenStudy (anonymous):

^^ the attachment are the answer choices

OpenStudy (anonymous):

@pooja195 we need you

OpenStudy (jhannybean):

is your function: \(\dfrac{\sqrt{5-x}}{\sqrt{x+5}}\) ?

OpenStudy (anonymous):

yes

OpenStudy (jhannybean):

\[\frac{\sqrt{5-x}\cdot \sqrt{x+5}}{\sqrt{x+5} \cdot \sqrt{x+5}}\]\[=\frac{\sqrt{25-x^2}}{x+5}\]

OpenStudy (jhannybean):

Im not too sure what you're looking for. Maybe you can specify your question. Are we looking for where x has a vertical asymptote?

OpenStudy (anonymous):

idk i just have the questiona dn the answers

OpenStudy (anonymous):

Here is the whole thing

OpenStudy (anonymous):

@Jhannybean

OpenStudy (anonymous):

@ganeshie8

OpenStudy (anonymous):

@Jhannybean @phi @Nnesha @ganeshie8 plz help

OpenStudy (millsemily):

Hold on

OpenStudy (anonymous):

@millsemily ok @ganeshie8 help?

OpenStudy (phi):

the domain are the "x" values you can use in your function and get a valid answer out. You can run into trouble if you divide by 0 (that is not allowed) or taking the square root of a negative number (that is not allowed, if we want real numbers for the answer) Normally, we would figure out what x values will cause trouble, and "not allow them" what is left over is the domain.

OpenStudy (phi):

First, look at the bottom \[ \sqrt{x+5} \] if that is 0, we would be dividing by 0 , and we don't allow that. notice if x=-5 we get sqr(0)= 0. we x=-5 is not allowed

OpenStudy (phi):

still looking at the bottom: \[ \sqrt{x+5} \] if x is less than -5, for example -6 we would have \[ \sqrt{-6+5}= \sqrt{-1} \] and we do not allow taking the square root of a negative number so we now know \( x\le -5\) is not allowed what is allowed is \( x \gt -5\) next, look at the top \[ \sqrt{5-x}\] if x is bigger than 5, we will get a negative number inside the square root, and that is not allowed. so x must be less than or equal to 5 we have \[ -5 \lt x \le 5\]

OpenStudy (anonymous):

ok thanls so much @Nnesha

Nnesha (nnesha):

what ??? o.O

OpenStudy (anonymous):

For helping me understand the problem thanks @Nnesha